The standard molar free energies of formation of NO2(g) and N2O4(g) at 25C are 51.840 and 98.280kJ/mol, respectively. What is the Value of Kp for the following equilibrium at 25C?
Answer #1

2NO2<---> N2O4(g)

free energy change , deltaG= standard free energy of N2O4-2*standard free energy of NO2 = 98.280-2*51.840 = -5.4 Kj

1 and 2 are stoichiometric coefficients of N2O4 and NO2 respectively.

deltaG=-RT ln K

lnK= equilibrium constant = -delatG/RT= 5.4*1000/ (8.314*298)=2.18

K= 8.84

for the reverse reaction N2O4<----> 2NO2 , KP= 1/8.84= 0.1132